a person invested Rs.1000 at 5% interest rate and allow to grow till it increased by 80%. in how many years did this happen?
Answer by Theo(13342)

the increase is 80% of 1000 which is equal to 800.
the future value is 1000 + 800 = 1800.
the formula is f = p * (1+r)^n
f is the future value.
p is the present value.
r is the interest rate per time period.
n is the number of time periods.
the formula becomes:
1800 = 1000 * (1.05)^n
divide both sides by 1000 to get:
1.8 = 1.05^n
take the log of both sides to get:
log(1.8) = log(1.05^n)
this becomes:
log(1.8) = n * log(1.05)
divide both sides by log(1.05) to get:
log(1.8)/log(1.05) = n
solve for n to get n = log(1.8)/log(1.05) = 12.04723687
the 1000 will grow to 1800 in 12.04723687 years.
f = 1000 * (1.05)^12.04723687 which is equal to 1800.
solution is confirmed to be correct.